a+b=180 ,a-2b=30 , then find the value of a and b
Answer by algebrahouse.com(1659)

a + b = 180 ------> -a - b = -180 {multiplied top equation by -1}
a - 2b = 30

-a - b = -180 {new top equation}
a - 2b = 30 {original bottom equation}
-------------
-3b = -150 {added the two equations together}
b = 50 {divided both sides by -3}

a + b = 180 {original top equation}
a + 50 = 180 {substituted 50, in for b, into top equation}
a = 130 {subtracted 50 from both sides}

a = 130 and b = 50

For more help from me, visit: www.algebrahouse.com